A firefighting crew uses a water cannon that shoots water at 25m/s at a fixed angle of 53 degrees above the horizontal. The firefighters want to direct the water at a blaze that is 10.0m above ground level. How far from the building should they position their cannon? There are two possibilities; can you get them both?


1
Expert's answer
2021-06-07T09:32:33-0400

Projectile motion relations:

"y-y_0=v_{0y}t+ \frac{1}{2}a_yt^2 \;(1)\\

x-x_0=x_{0x}t+ \frac{1}{2}a_xt^2 \;(2)"

where x,y are positions on x/y axis, "x_0, \;y_0" are initial positions, "v_{0x}, \;v_{0y}" are initial velocities and "a_x, \;a_y" are accelerations. We can choose that the positive direction of x-axis is to the right and the positive y-direction is upwards. Problem asks us to find "x-x_0\u2261 D" distance that we can position our water cannon if we know height H of the spot that we want to hit "(y-y_0\u2261H)" , the initial speed of the cannon, and angle of the water. So, first, we can use equation (1) to find time t

"H=v_0sinα_0t - \frac{1}{2}gt^2 \\

0=-4.9t^2 + 25 \times sin 53° \times t -10 \\

t_1=3.49 \;s \\

t_2=0.58 \;s"

If we put results into equation (2), we can get distances that we can place our water gun:

"D_1=v_0cosα_0t_1 \\

=25 \times cos53° \times 3.49 = 52.5 \;m \\

D_2 = v_0cosα_0t_2 \\

=25 \times cos53° \times 0.58 = 8.7 \;m"

Water cannon can be placed at distances 52.5 m and 8.7 m.